The chance
that a U.S. police chief believes the death penalty
“significantly reduces the number of homicides” is
1 in 4. If a random sample of 8 police chiefs is selected,
find the probability that at most 3 believe that the
death penalty significantly reduces the number of
homicides

Answer by Boreal(15235)

look at 0,1,2,3
for 0 it is 0.75^8=0.1001
for 1 it is 8*0.75^7*0.25=0.2670
for 2 it is 8C2*0.75^6*0.25^2=0.3115
for 3 it is 8C3*0.75^5*0.25^3=0.2076
add them up and the answer is 0.8862
or 2nd VARS binomcdf (8,0.25,3) (it does binomial from left end). and get same answer.